There are 4 into 10 power 24 gas molecules in a vessel at 50k temperature the pressure of the gas in the vessel is 0.03 atmospheric calculate the volume of the vessel

Grade:11

1 Answers

Hastin
5 years ago
IN,Ideal gas eq PV=nRT,R=0.0821(withr respect to given dimensions),T=50K,n=N/NA=4*1023/6*1023=2/3 and P=0.03atm (given).So put the values and get answer